A corporation's pension plan provides a lifetime annual income to its employees upon retirement at age 65. The plan provides 2% for each year of service of the employee's salary upon retirement. Those who retire before 65 have their benefit reduced by 1.6% for each year before 65 that they retire. Russ retires at age 63 with 23 years of service. If his salary upon retirement is $72928, what is his annual pension benefit? Round your answer to the nearest dollar.

I Answered: 6709, but the Correct Answer is 32473.0. Can someone show me how to do it correctly?
Answer #1

First we compute the benefit without deduction = Salary *2% * Years of service

= 72928*2%*23

= 33546.88

Deduction % = 1.6%*(65-63) = 3.2%

Hence the pension benefit =  benefit without deduction* (1- Deduction %)

= 33546.88*(1-0.032)

= $32473.38